de ph 10 300 (S, 30) ro to
Answer #1

Answer we can draw Rodo, Plot 忒 t-e Ge) H s) Gokn=01(s), H(s) OPen loop here G(s) i SoS十 hove G(s- 155 1300log to Lake o ot Lo 10 Mi 40 dB Since hore i one pole at ortin,t slope is -2o d BSlope of the Rondtont o ranae ec ec 20 dec 0-1 3 Exact plot D。 -45。 to Symptottc Plot61(s) 1 +G(S) s(s+30) 50S+ 300 503+ 300 s (st30 I 50s+ 300 s (S+30)+ ISOSt300 503+30O S+ 30st 150s+-306 S+ 180s +300= S, 180s+ 300 1o 10 5.16